OpenStudy (anonymous):

Center(9,−4) with a radius of 3 units anyone?

OpenStudy (anonymous):

you want the equation for the circle?

OpenStudy (anonymous):

general form is \[(x-h)^2+(y-k)^2=r^2\] here \(h=9,k=-4,r=3\)

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

i got this but it was wrong(x-9)^2+(y-(-4)^2=9

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

probably did not like the \(--4\) part

OpenStudy (anonymous):

try \[(x-9)^2+(y+4)^2=9\]